Question 12 345678 11 12 13 Assignment Submission Description Chapters 9 and 10. Momentum and the start of rotation For this assgrmert, you ent a by question parts. The number of submissions nemaining for each question part only changes you submit or change the answer Assignment Scoring Your last submission is used for your score 0. -2 points SerPSES 10 P.005 W y Notes Ask Your A wheel starts from rest and rotates with constant angular acceleration to reach an angular speed of 11.4 rad/s in 2.96 s (a) Find the magnitude of the angular acceleration of the wheel rad/s2 (b) Find the angle in radians through which it rotates in this time interval. rad Need Help? ǐ.linen.I Lentru Submit Answer Save ProgressPractice A Another Version << View Previous Question Question 10 of 13 View Next Question>> Home My Assignments Extension Request
Answer #1

Initial angular speed of the wheel, \small \omega0 = 0 rad / sec (since wheel starts from rest)

Final angular speed of the wheel, \small \omega = 11.4 rad / sec

Let \small \alpha is the angular acceleration of the wheel

Using Newton's angular motion equation:

\small \omega = \small \omega0 + \small \alpha t

here, t is the time in sec., t = 2.96 sec.

\small \Rightarrow 11.4 = 0 +  \small \alpha * 2.96  

\small \Rightarrow \alpha = 11.4 / 2.96

  \small \alpha = 3.85 rad. / sec.2

b)

Using Newton's angular motion equation:

  \small \theta = \omega0 t + 1/2 * \small \alphat2

here \small \theta is angle of rotation

\small \Rightarrow \theta = 0 * 2.96 +( 1/2)* 3.85 * 2.962

\small \theta = 16.87 rad